# Bismut torsion parallel metrics with constant holomorphic sectional curvature

### v1

@@ Abstract.
Zheng is the corresponding author. He is partially supported by National Natural Science Foundations of China with the grant No.12071050 and 12141101, Chongqing grant cstc2021ycjh-bgzxm0139, Chongqing Normal University grant 19XRC001, and is supported by the 111 Project D21024.
@@ 1. Introduction and statement of results
In dimension $2$, BTP = BKL = Vaisman. Such surfaces were fully classified by Belgun [5] in 2000. But when $n\geq 3$, BKL and Vaisman are disjoint, and their union is a proper subset of the set of all non-balanced BTP manifolds. Also, when $n\geq 3$, there are BTP manifolds that are balanced (and non-Kähler). Balanced BTP manifolds form a highly restrictive and interesting set. It contains examples that are Chern flat or Fano. We refer the readers to the preprint [26] and the references therein for more discussions on BTP manifolds.
The main result of this note is the following
@@
Consider compact Bismut torsion-parallel (BTP) manifolds. For all non-balanced ones and all three-dimensional balanced ones, Conjectures 1 and 2 hold.
Since Vaisman manifolds are non-balanced BTP, an immediate consequence to Theorem 1 is the following:
@@
Let $(M^{n},g)$ be a compact Hermitian manifold with $n\geq 2$. Assume that Bismut holomorphic sectional curvature $H^{b}=c$ is a non-zero constant. Then $g$ must be Kähler.
@@
Note that here we omitted the $c=0$ case. As proved in [10], for $n=2$ the statement $H^{b}=0\,\Longrightarrow\,R^{b}=0$ is valid, but for $n\geq 3$, there are examples of compact Hermitian manifolds with $H^{b}=0$ but with $R^{b}\not\equiv 0$. In [10, Lemma 5] it was shown that any isosceles Hopf manifold equipped with the standard metric will have $H^{b}=0$. But for $n\geq 3$, it does not have vanishing $R^{b}$. These ‘counterexamples’ on the other hand seem to be rather restrictive, so in [10] we also asked the following question:
@@
The reason to exclude the $t=-1$ case is that, as in the Bismut case, there are compact Hermitian manifolds with $H^{(-1)}=0$ but $R^{(-1)}\not\equiv 0$ when $n\geq 3$. So for this connection, the question should be proposed just like the Bismut case. Therefore in Conjecture 3 and Question 3 we will include $\nabla^{(-1)}$ there.
@@ 2. Preliminaries
for any $2$-plane $\pi$ spanned by $x$ and $y$ in the tangent space $T_{p}M$. Here we wrote $g=\langle\,,\rangle$ and as usual $|x\wedge y|^{2}=|x|^{2}|y|^{2}-\langle x,y\rangle^{2}$. Clearly, this value is independent of the choice of the basis $\{x,y\}$ of $\pi$. Since our metric $g$ is Hermitian, namely, $\langle Jx,Jy\rangle=\langle x,y\rangle$ for any $x,y$, we can define the holomorphic sectional curvature $H^{D}$ of $D$ as the restriction of $K^{D}$ on all $J$-invariant $2$-planes:
@@
[…]
Let $e=\{e_{1},\ldots,e_{n}\}$ be a local unitary frame, namely, each $e_{i}$ is a complex tangent vector field of type $(1,0)$ and $\langle e_{i},\overline{e}_{j}\rangle=\delta_{ij}$ for $1\leq i,j\leq n$. Denote by $\nabla^{r}$, $\nabla$, and $\nabla^{b}$ the Riemannian (Levi-Civita), Chern, and Bismut connection of $(M^{n},g)$, by $T$, $T^{b}$ the torsion of $\nabla$ and $\nabla^{b}$, and by $R^{r}$, $R$, $R^{b}$ the corresponding curvature tensor. Under the frame $e$, write
@@
¯$\overline{\partial}\varphi_{r}=-\sum_{i,j}\overline{T^{i}_{jr}}\,\varphi_{i}\wedge\overline{\varphi}_{j}\ \ \mbox{at}\ p.$
@@
¯$\overline{\partial}\gamma^{\prime}_{k\ell}=\overline{\partial}(T^{\ell}_{kr}\varphi_{r})=T^{\ell}_{kr,\bar{j}}\overline{\varphi}_{j}\wedge\varphi_{r}-T^{\ell}_{kr}\overline{T^{i}_{jr}}\varphi_{i}\wedge\overline{\varphi}_{j},$
@@ 3. BTP manifolds with constant holomorphic sectional curvature
[…]
@@
$H^{(t)}=c$$\Longleftrightarrow$ $\ \widehat{R^{(t)}}_{i\bar{j}k\bar{\ell}}=\frac{c}{2}(\delta_{ij}\delta_{k\ell}+\delta_{i\ell}\delta_{kj}),$
@@
Letting $i=n$ we deduce $c=0$, therefore $0=\frac{t^{2}-1}{4}|a_{i}|^{2}$. If $t\neq\pm 1$, then again it leads to $a_{i}=0$ for all $i$ and we get a contradiction. So for any non-balanced BTP manifolds and for any $t\neq\pm 1$, the $t$-Gauduchon holomorphic sectional curvature $H^{(t)}$ cannot be a constant. Note that when $t=0$ it is just the Chern connection case. For the $t=1$ or $-1$ cases, by letting $i=j=k=\ell=n$ in the formula in either Lemma 7 or Lemma 8, we end up with $c=0$. So for non-balanced BTP manifolds, the holomorphic sectional curvature of $\nabla^{b}$ or $\nabla^{(-1)}$ cannot be a non-zero constant. This completes the proof of Theorems 1 and 4 in the case when the BTP manifold is assumed to be non-balanced. ∎
@@
    (10) $\Theta^{b}=\left[\begin{array}[]{ccc}\alpha+\beta&0&0\\ 0&\alpha&\sigma\\ 0&-\bar{\sigma}&\beta\end{array}\right],\ \ \ \ \ \ \ \ \ \left\{\begin{array}[]{lll}\alpha=\varphi_{1\bar{1}}+(1\!-\!b)\varphi_{2\bar{2}}+b\varphi_{3\bar{3}}+t\varphi_{2\bar{3}}+\bar{t}\varphi_{3\bar{2}},&&\\ \beta=\varphi_{1\bar{1}}+b\varphi_{2\bar{2}}+(1\!-\!b)\varphi_{3\bar{3}}-t\varphi_{2\bar{3}}-\bar{t}\varphi_{3\bar{2}},&&\\ \sigma=t\varphi_{2\bar{2}}-t\varphi_{3\bar{3}}+s\varphi_{2\bar{3}}+(1\!+\!b)\varphi_{3\bar{2}},&&\end{array}\right.$
@@
    (11) $\Theta^{b}=\left[\begin{array}[]{ccc}d\alpha&d\beta_{0}&\\ -d\beta_{0}&d\alpha&\\ &&0\end{array}\right],\ \ \ \ \ \ \ \ \ \left\{\begin{array}[]{ll}d\alpha=\ x(\varphi_{1\bar{1}}+\varphi_{2\bar{2}})+iy(\varphi_{2\bar{1}}-\varphi_{1\bar{2}}),&\\ d\beta_{0}=-iy(\varphi_{1\bar{1}}+\varphi_{2\bar{2}})+(x\!-\!2)(\varphi_{2\bar{1}}-\varphi_{1\bar{2}}),&\end{array}\right.$
@@
Note that here we used the fact that $T^{i}_{i\ast}=0$ since $e$ is special. When $i\neq k$, the sigma part in the right hand side of the above equation becomes $-\frac{1}{2}a_{j}^{2}+\frac{3}{8}(a_{k}^{2}+a_{i}^{2})$, where $j\neq i,k$. Therefore we have
(12) $\left\{\begin{array}[]{ll}R^{b}_{i\bar{i}i\bar{i}}=c,&\\ R^{b}_{i\bar{i}k\bar{k}}=\frac{c}{2}-\frac{1}{2}a_{j}^{2}+\frac{3}{8}(a_{k}^{2}+a_{i}^{2}),\ \ \ \mbox{if}\ \{i,j,k\}=\{1,2,3\}&\end{array}\right.$
@@
When $i=k$, it is clearly equal to zero, while when $i\neq k$, it equals to $-|a_{j}|^{2}+|a_{k}|^{2}=0$, where $\{i,j,k\}=\{1,2,3\}$. That is, $R^{b}_{i\bar{i}k\bar{k}}=0$ for all $1\leq i,k\leq 3$. Compare with (12), we get $c=0$ and $0=\frac{1}{4}a_{1}^{2}$, which is a contradiction.
@@
(13) $\left\{\begin{array}[]{ll}R^{b}_{i\bar{i}i\bar{i}}=c,&\\ R^{b}_{i\bar{i}k\bar{k}}=\frac{c}{2}-\frac{1}{2}a_{j}^{2}+\frac{t^{2}+1}{4}(a_{k}^{2}+a_{i}^{2}),\ \ \ \mbox{if}\ \{i,j,k\}=\{1,2,3\}.&\end{array}\right.$
@@
For $t=1$, namely the Bismut connection case, by (11) we have $R^{b}_{3\bar{3}3\bar{3}}=0$ and $R^{b}_{1\bar{1}1\bar{1}}=x$. So if $H^{b}$ is a constant, then the constant must be 00 and $x=0$. Given any tangent direction $X=\sum_{i}x_{i}e_{i}$, write $p=|x_{1}|^{2}+|x_{2}|^{2}$, $q=x_{1}\overline{x}_{2}-x_{2}\overline{x}_{1}$. By (11) we have
@@
Clearly, for any fixed $y$ value, this expression cannot be identically zero for all choices of $X$. That is, $H^{b}$ can never be a constant. Similarly, for the ‘mirror’ connection $\nabla^{(-1)}$, by letting $t=-1$ in Lemma 4 we get
@@
As we have noted in the proofs, for BTP manifold, the $t=-1$ case of Lemma 4 gives us
@@
Acknowledgments. The second named author would like to thank Haojie Chen, Xiaolan Nie, Kai Tang, Bo Yang, Yashan Zhang, and Quanting Zhao for their interests or helpful discussions.

### v2

@@ Abstract.
Zheng is the corresponding author. He is partially supported by National Natural Science Foundations of China with the grant No.12071050 and 12141101, and is supported by the 111 Project D21024.
@@ 1. Introduction and statement of results
In dimension $2$, BTP = BKL = Vaisman. Such surfaces were fully classified by Belgun [5] in 2000. But when $n\geq 3$, BKL and Vaisman are disjoint, and their union is a proper subset of the set of all non-balanced BTP manifolds. Also, when $n\geq 3$, there are BTP manifolds that are balanced (and non-Kähler). Balanced BTP manifolds form a highly restrictive and interesting set. It contains examples that are Chern flat or Fano. We refer the readers to the preprints [27, 28] (which are rewriting of the older file [26]) and the references therein for more discussions on BTP manifolds.
We will say that Conjecture 1 (or 2) holds for a particular class of Hermitian manifolds, if any manifold in that class cannot have constant Chern (or Riemannian) holomorphic sectional curvature unless it is Kähler (hence a complex space form) or it is Chern (or Riemannian) flat. The main result of this note is the following
@@
Conjectures 1 and 2 hold for all non-balanced BTP manifolds and all three-dimensional balanced BTP manifolds.
Note that by [22, Theorem 3], any compact Hermitian manifold that is Chern (or Riemannian) flat (or more generally, Chern Kähler-like or Riemannian Kähler-like) must be balanced. So for the class ‘non-balanced BTP manifolds’ in Theorem 1, it simply says that such a manifold cannot have constant Chern (or Riemannian) holomorphic sectional curvature. Also, since all Vaisman manifolds are non-balanced BTP, an immediate consequence to Theorem 1 is the following:
@@
Let $(M^{n},g)$ be a compact Hermitian manifold with $n\geq 2$. Assume that Bismut holomorphic sectional curvature $H^{b}$ is a non-zero constant. Then $g$ must be Kähler.
@@
Note that here we omitted the $H^{b}=0$ case. As proved in [10], for $n=2$ the statement $H^{b}=0\,\Longrightarrow\,R^{b}=0$" is valid, but for $n\geq 3$, there are examples of compact Hermitian manifolds with $H^{b}=0$ but with $R^{b}\not\equiv 0$. In [10, Lemma 5] it was shown that any isosceles Hopf manifold equipped with the standard metric will have $H^{b}=0$. But for $n\geq 3$, it does not have vanishing $R^{b}$. These ‘counterexamples’ on the other hand seem to be rather restrictive, so in [10] we also asked the following question:
@@
The reason to exclude the t=1t=-1 case is that, as in the Bismut case, there are compact Hermitian manifolds with $H^{(-1)}=0$ but $R^{(-1)}\not\equiv 0$ when $n\geq 3$. So for this connection, the question should be proposed just like the Bismut case. Therefore in Conjecture 3 and Question 3 we will include $\nabla^{(-1)}$ there.
@@ 2. Preliminaries
for any $2$-plane $\pi$ spanned by $x$ and $y$ in the tangent space $T_{p}M$. Here we wrote $g=\langle\,$$\rangle$ and as usual $|x\wedge y|^{2}=|x|^{2}|y|^{2}-\langle x,y\rangle^{2}$. Clearly, this value is independent of the choice of the basis $\{x,y\}$ of $\pi$. Since our metric $g$ is Hermitian, namely, $\langle Jx,Jy\rangle=\langle x,y\rangle$ for any $x,y$, we can define the holomorphic sectional curvature $H^{D}$ of $D$ as the restriction of $K^{D}$ on all $J$-invariant 22-planes:
@@
where X=xiJxX=x-iJx is the complex tangent vector of type $(1,0)$ corresponding to $x$. Here and below we have extended $\langle\,$$\rangle$ and $R^{D}$ complex multi-linearly.
Let $e=\{e_{1},\ldots,e_{n}\}$ be a local unitary frame, namely, each $e_{i}$ is a complex tangent vector field of type $(1,0)$ and $\langle e_{i},\overline{e}_{j}\rangle=\delta_{ij}$ for $1\leq i,j\leq n$. Denote by $\nabla^{r}$, $\nabla$, and $\nabla^{b}$ the Riemannian (Levi-Civita), Chern, and Bismut connection of $(M^{n},g)$, by $T$, $T^{b}$ the torsion of $\nabla$ and $\nabla^{b}$, and by $^{r}, RR$, $R^{b}$ the corresponding curvature tensor. Under the frame $e$, write
@@
$\overline{\partial}\varphi_{r}=-\sum_{i,j}\overline{T^{i}_{jr}}\,\varphi_{i}\wedge\overline{\varphi}_{j}\ \ \mbox{at}\ p.$
@@
$\overline{\partial}\gamma^{\prime}_{k\ell}=\overline{\partial}(T^{\ell}_{kr}\varphi_{r})=T^{\ell}_{kr,\bar{j}}\overline{\varphi}_{j}\wedge\varphi_{r}-T^{\ell}_{kr}\overline{T^{i}_{jr}}\varphi_{i}\wedge\overline{\varphi}_{j},$
@@ 3. BTP manifolds with constant holomorphic sectional curvature
$\widehat{P}_{i\bar{j}k\bar{\ell}}=\frac{c}{2}(\delta_{ij}\delta_{k\ell}+\delta_{i\ell}\delta_{kj}),\ \ \ \ \ \forall\ 1\leq i,j,k,\ell\leq n.$
@@
$H^{(t)}=c\ \displaystyle\Longleftrightarrow$ $\ \widehat{R^{(t)}}_{i\bar{j}k\bar{\ell}}=\frac{c}{2}(\delta_{ij}\delta_{k\ell}+\delta_{i\ell}\delta_{kj}),$
@@
Letting $i=n$ we deduce $c=0$, therefore $0=\frac{t^{2}-1}{4}|a_{i}|^{2}$. If $t\neq\pm 1$, then again it leads to $a_{i}=0$ for all $i$ and we get a contradiction. So for any non-balanced BTP manifolds and for any $t\neq\pm 1$, the $t$-Gauduchon holomorphic sectional curvature $H^{(t)}$ cannot be a constant. Note that when $t=0$ it is just the Chern connection case. For the $t=1$ or 1-1 cases, by letting $i=j=k=\ell=n$ in the formula in either Lemma 7 or Lemma 8, we end up with $c=0$. So for non-balanced BTP manifolds, the holomorphic sectional curvature of $\nabla^{b}$ or $\nabla^{(-1)}$ cannot be a non-zero constant. This completes the proof of Theorems 1 and 4 in the case when the BTP manifold is assumed to be non-balanced. ∎
@@
[…]
@@
[…]
@@
Note that here we used the fact that $T^{i}_{i\ast}=0$ since $e$ is special. When $i\neq k$, the sigma part in the right hand side of the above equation becomes $\frac{1}{2}a_{j}^{2}+\frac{3}{8}(a_{k}^{2}+a_{i}^{2})$, where $j\neq i,k$. Therefore we have
[…]
@@
When $i=k$, it is clearly equal to zero, while when $i\neq k$, it equals to $|a_{j}|^{2}+|a_{k}|^{2}=0$, where $\{i,j,k\}=\{1,2,3\}$. That is, $R^{b}_{i\bar{i}k\bar{k}}=0$ for all $1\leq i,k\leq 3$. Compare with (12), we get $c=0$ and $0=\frac{1}{4}a_{1}^{2}$, which is a contradiction.
@@
[…]
@@
For $t=1$, namely the Bismut connection case, by (11) we have $R^{b}_{3\bar{3}3\bar{3}}=0$ and $R^{b}_{1\bar{1}1\bar{1}}=x$. So if $H^{b}$ is a constant, then the constant must be 0 and $x=0$. Given any tangent direction $X=\sum_{i}x_{i}e_{i}$, write $p=|x_{1}|^{2}+|x_{2}|^{2}$, $q=x_{1}\overline{x}_{2}-x_{2}\overline{x}_{1}$. By (11) we have
@@
Clearly, for any fixed $y$ value, this expression cannot be identically zero for all choices of $X$. That is, $H^{b}$ can never be a constant. Similarly, for the ‘mirror’ connection $\nabla^{(-1)}$, by letting t=1t=-1 in Lemma 4 we get
@@
As we have noted in the proofs, for BTP manifold, the t=1t=-1 case of Lemma 4 gives us
@@
Acknowledgments. The second named author would like to thank Haojie Chen, Xiaolan Nie, Kai Tang, Bo Yang, Yashan Zhang, and Quanting Zhao for their interest and helpful discussion.